# Dynamical cluster-based optimization of tensor network algorithms for quantum circuit simulations

### v1

@@ Appendix A Gates used in random-structure quantum circuits
$4]{X}=e^{i\frac{\pi}{8}}\begin{pmatrix}\cos\left(\frac{\pi}{8}\right)&-i\sin\left(\frac{\pi}{8}\right)\\ -i\sin\left(\frac{\pi}{8}\right)&\cos\left(\frac{\pi}{8}\right)\end{pmatrix},$ (18)
@@
Appendix B Further results for DMRG: fixed vs adaptive grouping
Figure 12: Average speedup of the DMRG algorithm with adaptive grouping against fixed grouping for 20 realizations of non-Clifford random structure quantum circuits of various sizes.
In this Appendix, we present additional results from the DMRG benchmarks illustrated in Sec. IV.3. In particular, our objective is to test our adaptive grouping scheme by comparing its runtime with a fixed grouping approach. In the fixed grouping implementation, we relax the constraints of the algorithm by determining the grouping based on the gate distribution across the entire circuit prior to the first iteration. This predetermined grouping is then employed across all subsequent DMRG iterations. A key difference with this approach is that, as the grouping is fixed throughout the algorithm, no decomposition steps are required at the beginning of each new iteration.
@@ Appendix B Further results for DMRG: fixed vs adaptive grouping
where $\text{runtime}_{\text{fixed}}(Q,L)$ is the runtime for DMRG with fixed grouping for a circuit of $Q$ qubits and $L$ layers, and $\text{runtime}_{\text{adaptive}}(Q,L)$ is the runtime with adaptive grouping. As shown in the figure, all points lie above $1.0$, meaning that using the adaptive grouping scheme in the DMRG algorithm yields a lower runtime on average with respect to fixed grouping. Moreover, the speedup tends to increase for both large and small values of $L_{\mathrm{max}}$, whereas for intermediate values of $L_{\mathrm{max}}$ the speedup is very slight.

### v2

@@ Appendix A Gates used in random-structured quantum circuits
$\sqrt[4]{X}=e^{i\frac{\pi}{8}}\begin{pmatrix}\cos\left(\frac{\pi}{8}\right)&-i\sin\left(\frac{\pi}{8}\right)\\ -i\sin\left(\frac{\pi}{8}\right)&\cos\left(\frac{\pi}{8}\right)\end{pmatrix}$, (18)
@@
Appendix B Further results for DMRG: fixed versus adaptive grouping
Figure 12: Average speedup of the DMRG algorithm with adaptive grouping against fixed grouping for 20 realizations of non-Clifford random-structured quantum circuits of various sizes.
In this Appendix, we present additional results from the DMRG benchmarks illustrated in Sec. IV.3. In particular, our objective is to test our dynamical adaptive grouping routine by comparing its runtime with a fixed grouping approach. In the fixed grouping implementation, we relax the constraints of the algorithm by determining the grouping based on the gate distribution across the entire circuit prior to the first iteration. This predetermined grouping is then employed across all subsequent DMRG iterations. A key difference with this approach is that, as the grouping is fixed throughout the algorithm, no decomposition steps are required at the beginning of each new iteration.
@@ Appendix B Further results for DMRG: fixed versus adaptive grouping
where $\text{runtime}_{\text{fixed}}(Q,L)$ is the runtime for DMRG with fixed grouping for a circuit of $Q$ qubits and $L$ layers, and $\text{runtime}_{\text{adaptive}}(Q,L)$ is the runtime with adaptive grouping. As shown in the figure, all points lie above $1.0$, meaning that using the dynamical adaptive grouping routine in the DMRG algorithm yields a lower runtime on average with respect to fixed grouping. Moreover, the speedup tends to increase for both large and small values of $L_{\mathrm{max}}$, whereas for intermediate values of $L_{\mathrm{max}}$ the speedup is very slight.